Ann is 11 years ago. Her Aunt, Kate turned 47 today. After how many years will Kate be as old as Ann
Answer by ikleyn(52781)

.
Ann is 11 years ago. Her Aunt, Kate turned 47 today. After how many years will Kate be as old as Ann
~~~~~~~~~~~~~~


What this sentence of your post means ?

    " Ann is 11 years ago. "


Write clearly, P L E A S E.


///////////////


Comment from student : Ann is 11 years old. Her Aunt, Kate turned 47 today.
After how many years will Kate be 3 times as old as Ann.



My SOLUTION

Let it will happen in x years


    3*(11+x) = 47 + x


Simplify and find x

    33 + 3x = 47 + x

    3x - x = 47 - 33

       2x  =    14

        x  =    14/2 = 7.


ANSWER.  In 7 years.

Solved.